OpenStudy (anonymous):

what minimum concentration of Na2SO4 is required to cause the precipitation of the cation that precipitates first. (The cation is Ba+). consider that the solution is 1.1*10^-2 M Ba2+ and 1.8*10^-2 M Ca2+

OpenStudy (kira_yamato):

Do you have the Ksp of BaSO4 and CaSO4?

OpenStudy (kira_yamato):

\[\left[ Na_2SO_4 \right] = \left[ SO_4^{2-} \right] = \frac{K_{sp}(BaSO_4)}{[Ba^{2+}]}\]